OpenStudy (anonymous):

Find the solution of this system of equations. separate the x-and y- values with a comma. x=5+y 28x-9y=-12

OpenStudy (anonymous):

I will try to solve your system of equations. x=y+5;28x−9y=−12 Step: Solve x=y+5 for x: Step: Substitute (y+5) for x in 28x−9y=−12: 28x−9y=−12 28(y+5)−9y=−12 19y+140=−12(Simplify both sides of the equation) 19y+140+−140=−12+−140(Add -140 to both sides) 19y=−152 19y19=−15219(Divide both sides by 19) y=−8 Step: Substitute (−8) for y in x=y+5: x=y+5 x=−8+5 x=−3(Simplify both sides of the equation) Answer: x=−3 and y=−8

OpenStudy (sachintha):

notgoodatmath16 did you get it?

OpenStudy (anonymous):

Yes thank you(: